OpenStudy (anonymous):

Note:Base 4= B4 logB4(x+3)+logB4(x-3)=2

OpenStudy (anonymous):

step is to convert it to log4[(x-3)(x+3)] = 2 Now we raise 4 to the power of both sides (x-3)(x+3) = 4^2 Simplify x^2 - 9 = 16 Add 9 to both sides x^2 = 25 x = 5 or x = -5

OpenStudy (mathmale):

Shay, I'm so sorry about @Raymond67 's totally unacceptable comment.

OpenStudy (anonymous):

@LegoMyEgo so you change it to exponential form?

OpenStudy (anonymous):

@mathmale Thanks for removing it.

OpenStudy (mathmale):

Several things: First, it's true that logB4(x+3)+logB4(x-3) = logB4 [(x+3)(x-3)]. This stems from one of the most important rules of logs.

OpenStudy (anonymous):

how do you know when to change it?

OpenStudy (mathmale):

Secondly, y=logB4 x and y=4^x are INVERSE FUNCTIONS.

OpenStudy (mathmale):

Are you, Shay, familiar with this term, inverse function?

OpenStudy (anonymous):

Shay, do you wan Mathmale to explain it to you?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

kk

OpenStudy (mathmale):

Shay: Rule 1: log a + log b = log a*b (multiplication rule) Rule 2: log a - log b = log (a/b) (division rule) Rule 3: log a^b = b*log a (power rule) I'd be happy to discuss these with you further if you'd like. But back to the original logarithmic equation:

OpenStudy (mathmale):

Using Rule 1: logB4 (x+3) + logB4 (x-3) = logB4 (x+3)(x-3). Before we move on, are you OK with this?

OpenStudy (anonymous):

yes @ mathmale